Michael Jordan had a vertical jump of about 1.22m. Assuming no air resistance, at what speed did he leave the ground?


1
Expert's answer
2021-03-04T11:51:03-0500
"v^2=2gh\\v^2=2(9.8)(1.22)\\v=4.89\frac{m}{s}"